11 Isotonic and Buffer Solutions 193 CHLORIDE EQUIVALENTS (E-VALUES) (Continuedh Molecular Weight lons(E-valuer Sodium Chloride Equivalent 92 0.35 0.17 354 0.09 182 759 0.18 aH0 0.11 0.13 0.27 0.22 0.18 0.24 247 372 G potassium 0.32 perylephrine hydrochloride 413 245 74.5 166 101 136 273 331 438 170 0.16 0.24 0.76 salicylate 0.58 0.43 0.21 phosphate, monobasic 0.15 0.12 0.33 0.65 0.42 0.6 acaine hydrochloride hydrochloride Scopolamine hydrobromide 3H.O er nrate Sodium bicarbonate Sodum borate-10H,0 Sadum carbonate-H0 dum chloride Sdum citrate 2H.0 dum iodide dum lactate 84 381 124 0.31 0.39 0.52 0.53 150 112 142 268 120 138 301 481 idum phosphate, dibasic, anhydrous m phosphate, dibasic-7H,0 um phosphate, monobasic, anhydrous 0.29 0.49 0.42 0.18 phosphate, monobasic-H,0 ecane hydrochloride oine hydrochloride erline hydrochloride 0.25 0.07 432
media%2F52b%2F52b543cc-62b2-4a6b-aa1d-2d
Answer #1

Molecular weight of homatropine hydrobromide = 356 grams

1% homatropine hydrobromide is needed. 0.356 grams in 100 mL is considered 1% solution. So, 15 ml solution require, 0.534 grams

E value of homatropine hydrobromide is, 0.17

So, the volume of water = 0.534*0.17/ 0.009 = 1.008 or 1

The rest of the volume of water = 15-1 = 14 mL

The rest of the volume, i..e 14 mL is isotonic by boric acid.

Thus, the weight of boric acid needed is, = 14*0.009/0.53 = 0.23 grams (approximately)